### v3

@@ Abstract
Artificial intelligence (AI) has seen remarkable advancements across various domains, including natural language processing, computer vision, autonomous vehicles, and biology. However, the rapid expansion of AI technologies has escalated the demand for more powerful computing resources. As digital computing approaches fundamental limits, neuromorphic photonics emerges as a promising platform to complement existing digital systems. In neuromorphic photonic computing, photonic devices are controlled using analog signals. This necessitates the use of digital-to-analog converters (DAC) and analog-to-digital converters (ADC) for interfacing with these devices during inference and training. However, data movement between memory and these converters in conventional von Neumann computing architectures consumes energy. To address this, analog memory co-located with photonic computing devices is proposed. This approach aims to reduce the reliance on DACs and minimize data movement to enhance compute efficiency. This paper demonstrates a monolithically integrated neuromorphic photonic circuit with co-located capacitive analog memory and analyzes analog memory specifications for neuromorphic photonic computing using the MNIST dataset as a benchmark.
@@ 1.1 Motivation
Artificial intelligence (AI) has profoundly influenced society and technology with advancements ranging from ChatGPT’s human-like dialogues (https://chat.openai.com/) to AlphaFold’s protein structure predictions [1, 2, 3]. Historically, AI’s progression has paralleled advances in CMOS electronics, driven by Moore’s law and Dennard scaling [4]. While Moore’s law has slowed down, AI’s performance demands have surged, with computational needs doubling every two months [4]. In parallel, neuromorphic engineering has emerged as a field aiming to align AI algorithms to hardware architectures that mirror their inherently distributed nature, with demonstrated applications in particle physics, nonlinear programming, and signal processing [5]. Among these innovations, neuromorphic photonics promise high bandwidth and low latency, providing a complementary opportunity to extend the domain of AI [5, 6, 7, 8]. However, most photonic processors have been limited to stationary weights, which are trained offline on digital computers and remain unchanged during inference, thus restricting their application potential. One reason this limitation arises is because analog photonic processors require analog-to-digital (ADC) and digital-to-analog (DAC) converters (Figure 1b), which become critical bottlenecks that affect bandwidth and energy efficiency as data transitions between digital memory and analog computation [9, 5, 10, 11]. During training, backpropagation requires computations (gradient and error calculations) and memory accesses (weight updates and temporary data storage for calculations), meaning that any part of the pipeline that requires data converters to compute or access memory can be a bottleneck. This work focuses on data conversion bottlenecks from memory and showcases energy savings and minimised latency from co-locating analog memory with analog optical computational units.
@@
Figure 1: Motivation for electro-optic processors with dynamic electro-optic analog memory (DEOAM). a) Dynamic electro-optic analog memory (DEOAM) consists of a capacitor connected to a PN junction microring resonator (MRR). The capacitor holds data on the MRR, thereby enacting a weight in the optical domain. Using DEOAM means DACs can be reused among rows of analog memory devices since the analog memory holds the signal on the MRR, which relaxes limitations on energy efficiency and bandwidth imposed by DACs. b) Conventional electro-optic processors require a dedicated SRAM-DAC for each PN junction MRR. The DAC needs to be constantly active to hold a signal on the MRR; therefore, implementing SRAM-DAC in a neuromorphic photonic processor requires a DAC for every MRR. c) If n represents the number of MRRs in the row and column of a neuromorphic photonic processor, scaling the conventional SRAM-DAC implementation incurs an n2 DAC penalty, whereas DEOAM scales by n. Write power and sampling rate for 32x32, 8-bit memory cell architectures are compared using each memory device and DACs from literature [12, 13, 14, 15, 16, 17, 18, 19, 20]. As DEOAM scales, DEOAM saves more power compared to SRAM-DAC.
@@
Analog memory can reduce the number of DACs in neural network architectures (Figure 1a). While optical analog memories (OAM) have been explored as in-memory computing devices, they currently face challenges, such as low yield and endurance, incompatibility with standard foundry processes, and complex fabrication processes [21, 22]. In contrast, analog electronic memories, such as dynamic random-access memory (DRAM), metal-oxide-metal capacitors, metal-insulator-metal capacitors, or metal-on-semiconductor capacitors, are foundry-compatible and included in conventional CMOS technologies. This compatibility makes integrating analog electronic memories into neuromorphic photonic hardware a more viable approach, potentially reducing the number of ADCs and DACs in applications, such as long-short term memory neural networks [23].
@@
Here, we propose a neuromorphic photonic processor with dynamic electro-optic analog memory (DEOAM, Figure 1a). This processor leverages the monolithic integration of nanophotonic and CMOS devices on the same silicon substrate. The proposed architecture (Figure 1a) merges the parallel information processing capabilities of photonics using wavelength-division multiplexing (WDM) with the low control complexity of crossbar arrays offered by electronics. As a proof-of-concept, we demonstrate neurosynaptic photonic weights, realized through a WDM array of tunable PN junction microring resonators (MRR) within a photonic weightbank and integrated with an array of capacitive DEOAM on the 90 nm GF9WG monolithic process. The MNIST dataset serves as a benchmark for training and inference and reveals tradeoffs in neural network design [24]. Our approach to integrating analog memory in neuromorphic photonic processors aims to create a flexible processor with high bandwidth, efficient training capabilities, and the ability to adapt through on-chip, online learning.
@@ 1.2 Dynamic Electro-Optic Analog Memory (DEOAM)
In a neuromorphic photonic processor with n rows and n columns of MRRs acting as weights, DEOAM reduces the number of DACs, thereby reducing power consumption. Since the conventional SRAM-DAC implementation requires each MRR to be paired with a DAC, scaling the implementation results in the number of DACs increasing by n2 (Figure 1b). The power consumed by SRAM-DAC is calculated as:
$P_{SRAM-DAC}=n^{2}P_{DAC}+P_{SRAM_{dynamic}}n+P_{SRAM_{static}}n^{2}$ (1)
@@
where PSRAM-DAC is the total SRAM-DAC array power consumption in W, and PSRAM and PDAC are the SRAM and DAC power consumption in W, respectively, which includes static and dynamic power consumption and accounts for varying bit precisions of the DAC. All surveyed DACs are demonstrated, and the SRAMs associated with each DAC are surveyed to be in the same process node [12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
In contrast, using DEOAM reduces the number of DACs below n2 (Figure 1a) because data is held directly on the photonic device via analog memory, eliminating the constant processing typically needed by DACs. This concept is similar to the functionality in modern liquid crystal active-matrix displays [25]. The power consumed by DEOAM is calculated as:
$P_{DEOAM}=(P_{DAC}+\alpha CfV^{2})n$ (2)
@@
where PDEOAM is the total DEOAM array power consumption, $\alpha$ is the activity factor (assumed to be 1), C is the load capacitance in F, f is the switching frequency in Hz and is half the sampling rate, and V is the switching voltage. DACs are still required in the DEOAM architecture to load initial weights from non-volatile, digital memory before beginning inference or training and to refresh analog memory during inference. However, this approach not only simplifies control complexities by reducing the number of electrical lines needed for DAC operations but also, with longer retention times, diminishes the need for DACs to continuously drive the photonic devices, leading to further power savings as the processor scales (Figure 1c, inset). Hence, integrating DEOAM and reducing the reliance on DACs paves the way for more efficient computing.
@@
In neuromorphic processors, analog memory is typically situated close to or directly integrated with the computing unit to minimize data movement, thereby reducing energy consumption and latency [26, 27]. Data movement in memory accesses can consume more than 10 times the energy of multiplication and addition operations [28]. Recent demonstrations of in-memory compute, such as Mythic’s Analog Matrix Processor or NeuRRAM, achieve up to tens of TOPS/W in compute efficiency by co-locating the analog memory with the compute unit [29, 30, 31, 32]. This highlights DEOAM’s close integration with the MRR, which acts as the compute unit in the optical domain, thereby minimizing data movement and enhancing compute efficiency.
@@ 2.1 Electro-Optic Characteristics
The neuromorphic photonic hardware with integrated DEOAM is shown in Figure 2. Key figures of merit for DEOAM are summarized in Table 1, and measurement results are shown in Figure 3. Figure 2a and b show the DEOAM circuit and its operation. Weights are loaded as voltages on the capacitive memory, affecting the MRR’s resonance. The capacitive memory holds the voltage on the MRR until the charge leaks away, returning the MRR to its resonance. Figure 2c shows the chip and measurement setup for static optical and time domain measurements. Free spectral range, inter-channel spacing, Q-factor, and extinction ratio are crucial optical characteristics that determine the maximum number of MRRs and weights in a weight bank in a specific wavelength range. Therefore, if additional weights are needed in each row of a weight bank, solutions like adding more cores, reusing hardware, or employing time multiplexing may be needed for processing neural network data. In a neuromorphic photonic weight bank, light modulation is initiated at the input data modulators. The light then travels through the weight banks before being absorbed by photodetectors (PD) and amplified by transimpedance amplifiers (TIA). The group index helps characterize the group delay, which determines the overall compute time when combined with the delay from the PD and TIA. Since the delay from the PD and TIA depends on their design and technology, we report compute time based primarily on the group delay. State-of-the-art demonstrations of PDs and TIAs yield latencies ranging in tens of ps [33, 34, 35, 36, 37]. The wavelength and modulation efficiencies are also important for characterizing the tunability and range of weights for the neural network.
@@
Figure 2: The architecture and operation of the neuromorphic photonic prototype with integrated DEOAM. a) The analog memory circuit consists of a 1x4 MRR weight bank with each MRR connected to a capacitive analog memory. Transmission gates are used to enable/disable (EN) writing and switches are used to reset (RST) the analog memories. Laser light is injected into grating couplers through the IN port and is measured at the THRU and DROP ports. b) The operation of the DEOAM circuit is shown. The PN junction MRR is modeled as a reversed biased diode, and the transmission gate is modeled as an ideal switch. At time 1, the laser has been tuned to the MRR’s resonance, and the transmission gate is connected such that the voltage source (VIN) can drive the analog memory. At time 2, voltage is driven onto the capacitor and MRR, causing the MRR’s resonance to shift. At time 3, the transmission gate disconnects the voltage source from the capacitor, and the capacitor holds the charge on the MRR until it leaks away, returning the MRR resonance back to its original state. c) The chip is wire bonded to a custom printed circuit board with thermal control and high speed electrical lines. Details about the hardware operation and measurement setup are expanded in Appendix A.
@@
Key FoMs to characterize the capacitive analog memory are retention time, write time, read time, and power consumption. Since data written to DEOAM is immediately available in the optical domain, read time is effectively zero. Retention time is dependent on leakage, and leakage is dependent on optical power and photoresponse of the PN junction MRR [38]. Figures 3a and b reveal that the retention time of 0.8345 ms for one time constant can be extended if lower optical power is used since lower leakage is observed, but this involves a tradeoff between retention time and signal-to-noise ratio. Energy consumption, estimated to be 55.97 pJ/write, is linked to capacitor leakage and charge replenishment to maintain weight values. Figures 3c, d, and e show that the write time is about 40 to 50 ns for one time constant, enabling MHz write speeds. From Figure 3, the analog memory bit precision is about 5 bits and can be extracted from the time domain responses based on $log_{2}(\frac{\overline{\mu}_{max}-\overline{\mu}_{min}}{\sigma})=N_{b}$ where $\overline{\mu}_{max}$ and $\overline{\mu}_{min}$ are the mean values at the maximum and minimum range of the analog memory, respectively, $\sigma$ is the standard deviation, and $N_{b}$ is the bit precision of the analog memory [39].
@@
Table 1: Experimental measurement results of the electro-optic analog memory circuit. Write time and retention time measured with 0.5 dBm optical bus power in the circuit. Details regarding measurement setup and optical characteristics can be found in Appendix A and B.
@@
Free Spectral Range FSR - 4.67 - nm
Full Width Half Max. FWHM - 0.019 - nm
@@
[…]
Bit Precision $N_{b}$ 5.33 5.65 5.97 bits
Capacitance CMEM - 14.005 - pF
[…]
Figure 3: Electro-optic characteristics of the DEOAM. a) Leakage shows a dependence on optical power incident on the doped MRR. Error bands represent the standard deviation in the leakage. b) Retention time from leakage is 0.835 ms for one time constant at 0.5 dBm optical bus power. The decay shows a nonlinear response from the ring’s Lorentzian response. c) With the transmission gate on, 2 V is written to analog memory using a 5 ns edge rate, and the resulting write time is 45.1 ns for one time constant. d) With 2 V on VIN and toggling the transmission gate from off to on, write time is 40.0 ns. e) With the transmission gate on and writing 0 V to analog memory at an edge rate of 5 ns, write time is 51.0 ns for one time constant. Further discussions on the electro-optic results are expanded in Appendix C.

### v5

@@ Abstract
[…]
@@ 1.1  Motivation
Artificial intelligence (AI) has profoundly influenced society and technology with advancements ranging from ChatGPT’s human-like dialogues (https://chat.openai.com/) to AlphaFold’s protein structure predictions [alquraishi_alphafold_2019, merolla_million_2014, bannon_computer_2019]. Historically, AI’s progression has paralleled advances in CMOS electronics, driven by Moore’s law and Dennard scaling [mehonic_brain-inspired_2022]. While Moore’s law has slowed down, AI’s performance demands have surged, with computational needs doubling every two months [mehonic_brain-inspired_2022]. In parallel, neuromorphic engineering has emerged as a field aiming to align AI algorithms to hardware architectures that mirror their inherently distributed nature, with demonstrated applications in particle physics, nonlinear programming, and signal processing [shastri_photonics_2021]. Among these innovations, neuromorphic photonics promise high bandwidth and low latency, providing a complementary opportunity to extend the domain of AI [shastri_photonics_2021, ferreira_de_lima_primer_2020, de_lima_machine_2019, tait_neuromorphic_2017]. The key advantages of neuromorphic photonic accelerators over purely electronic implementations include 1) massive parallelism using optical techniques, such as wavelength-division multiplexing (WDM), mode multiplexing (MDM), and time-domain multiplexing to carry multiple signals in a single interconnect and 2) lower energy-delay products in optical interconnects for multiply-accumulate (MAC) operations with potential speeds on the order of peta MACs per second per mm2 and energy efficiencies in the atto joule per MAC regime [feldmann_parallel_2021, khaled_fully_2024, lin_120_2024, ou_hypermultiplexed_2025, shastri_photonics_2021]. However, most photonic processors have been limited to stationary weights, which are trained offline on digital computers and remain unchanged during inference, thus restricting their application potential. One reason this limitation arises is because analog photonic processors require analog-to-digital (ADC) and digital-to-analog (DAC) converters (Figure 1b), which become critical bottlenecks that affect bandwidth and energy efficiency as data transitions between digital memory and analog computation [sun_full_2023, shastri_photonics_2021, antonik_online_2018, filipovich_silicon_2022]. During training, backpropagation requires computations (gradient and error calculations) and memory accesses (weight updates and temporary data storage for calculations), meaning that any part of the pipeline that requires data converters to compute or access memory can be a bottleneck. This work focuses on data conversion bottlenecks from memory and showcases energy savings and minimised latency from co-locating analog memory with analog optical computational units.
@@
Figure 1: Motivation for electro-optic processors with dynamic electro-optic analog memory (DEOAM). a) Dynamic electro-optic analog memory (DEOAM) consists of a capacitor connected to a PN junction microring resonator (MRR). The capacitor holds data on the MRR, thereby enacting a weight in the optical domain. Using DEOAM means DACs can be shared amongst columns of analog memory devices, updating them row by row since the analog memory holds the signal on the MRR, which relaxes limitations on energy efficiency and bandwidth imposed by DACs. b) Conventional electro-optic processors require a dedicated SRAM-DAC for each PN junction MRR. The DAC needs to be constantly active to hold a signal on the MRR; therefore, implementing SRAM-DAC in a neuromorphic photonic processor requires a DAC for every MRR. c) If $n$ represents the number of MRRs in the row and column of a neuromorphic photonic processor, scaling the conventional SRAM-DAC implementation incurs an $n^{2}$ DAC penalty, whereas DEOAM scales by $n$. Write power and sampling rate for 32x32, 8-bit memory cell architectures are compared using each memory device and DACs from literature [ma_12-bit_2015, ansari_5mw_2014, kumar_impact_2021, wu_130_2008, hong_28nm_2015, widmann_time-interleaved_2023, zhang_06-v_2018, kumar_performance_2017, rubino_880_2023]. As DEOAM scales, DEOAM saves more power compared to SRAM-DAC.
@@
Analog memory can reduce the number of DACs in neural network architectures (Figure 1a). While optical analog memories (OAM) have been explored as in-memory computing devices, they currently face challenges, such as low yield and endurance, incompatibility with standard foundry processes, and complex fabrication processes [youngblood_integrated_2023, tossoun_high-speed_2024]. Phase change materials (PCM) have been demonstrated in a photonic tensor core enabling 1012 MACs per second for inference tasks [feldmann_parallel_2021]. However, the endurance of PCMs can vary depending on the modulation mechanism (electrical, electrothermal, or optical), raising concerns about their suitability for neuromorphic training applications that typically require volatile memory [martin-monier_endurance_2022]. Optical memristors, which store memory by forming and dissolving filaments, offer minimal footprint, making them promising candidates for large scale integration. However, their limited endurance (approximately 2,000 write cycles) and challenges in compatibility with standard foundry processes remain key obstacles to broader adoption [liu_multilevel_2022, youngblood_integrated_2023]. In contrast, analog electronic memories, such as dynamic random-access memory (DRAM), metal-oxide-metal capacitors, metal-insulator-metal capacitors, or metal-on-semiconductor capacitors, are foundry-compatible and included in conventional CMOS technologies. This compatibility makes integrating analog electronic memories into neuromorphic photonic hardware a more viable approach, potentially reducing the number of ADCs and DACs in applications, such as long-short term memory neural networks [howard_photonic_2020].
@@
Here, we propose a neuromorphic photonic processor with dynamic electro-optic analog memory (DEOAM, Figure 1a). This processor leverages the monolithic integration of nanophotonic and CMOS devices on the same silicon substrate. The proposed architecture (Figure 1a) merges the parallel information processing capabilities of photonics using wavelength-division multiplexing (WDM) with the low control complexity of crossbar arrays offered by electronics. As a proof-of-concept, we demonstrate neurosynaptic photonic weights, realized through a WDM array of tunable PN junction microring resonators (MRR) within a photonic weightbank and integrated with an array of capacitive DEOAM on the 90 nm GF9WG monolithic process. The MNIST dataset serves as a benchmark for training and inference and reveals tradeoffs in neural network design [lecun_gradient-based_1998]. Our approach to integrating analog memory in neuromorphic photonic processors aims to create a flexible processor with high bandwidth, efficient training capabilities, and the ability to adapt through on-chip, online learning.
@@ 1.2  Dynamic Electro-Optic Analog Memory (DEOAM)
In a neuromorphic photonic processor with $n$ rows and $n$ columns of MRRs acting as weights, DEOAM reduces the number of DACs, thereby reducing power consumption. Since the conventional SRAM-DAC implementation requires each MRR to be paired with a DAC, scaling the implementation results in the number of DACs increasing by $n^{2}$ (Figure 1b). The power consumed by SRAM-DAC is calculated as:
$P_{\mathrm{SRAM-DAC}}=n^{2}(P_{\mathrm{DAC_{static}}}+P_{\mathrm{SRAM_{static}}})+n(P_{\mathrm{DAC_{dynamic}}}+P_{\mathrm{SRAM_{dynamic}}})$ (1)
@@
where $P_{\mathrm{SRAM-DAC}}$ is the total SRAM-DAC array power consumption in W, and $P_{\mathrm{SRAM}}$ and $P_{\mathrm{DAC}}$ are the SRAM and DAC power consumption in W, respectively, which includes static and dynamic power consumption and accounts for varying bit precisions of the DAC. All surveyed DACs are demonstrated, and the SRAMs associated with each DAC are surveyed to be in the same process node [ma_12-bit_2015, ansari_5mw_2014, kumar_impact_2021, wu_130_2008, hong_28nm_2015, widmann_time-interleaved_2023, zhang_06-v_2018, kumar_performance_2017, rubino_880_2023].
@@
In contrast, using DEOAM reduces the number of DACs below $n^{2}$ (Figure 1a) because data is held directly on the photonic device via analog memory, eliminating the constant processing typically needed by DACs. This concept is similar to the functionality in modern liquid crystal active-matrix displays [den_boer_active_2005]. The power consumed by DEOAM is calculated as:
$P_{\mathrm{DEOAM}}=n(P_{\mathrm{DAC_{static}}}+P_{\mathrm{DAC_{dynamic}}}+\alpha CfV^{2})$ (2)
@@
where $P_{\mathrm{DEOAM}}$ is the total DEOAM array power consumption, $\alpha$ is the activity factor (assumed to be 1), $C$ is the load capacitance in F, $f$ is the switching frequency in Hz and is half the sampling rate, and $V$ is the switching voltage. If we consider a memory element and DAC in both architectures, both architectures consume similar power when modulating the inputs (input to the SRAMs in the SRAM-DAC architecture and input to the DAC in the DEOAM architecture), but the main difference in power consumption is the SRAM to DAC power consumption in the SRAM-DAC architecture, the DAC to DEOAM power consumption in the DEOAM architecture, and the distribution of active DACs between the two architectures. In the SRAM-DAC architecture, power consumption begins when the SRAM changes state and ends when the DAC settles, meaning there is SRAM static and dynamic power and DAC static and dynamic power. In the DEOAM architecture, power consumption begins when the DAC changes state and ends when charge settles on the DEOAM, meaning there is DAC static and dynamic power and DEOAM dynamic power. Finally, there is the distribution of DACs from the architecture that determines the power consumption scaling. DACs are still required in the DEOAM architecture to load initial weights from non-volatile, digital memory before beginning inference or training and to refresh analog memory during inference. However, this approach not only simplifies control complexities by reducing the number of electrical lines needed for DAC operations but also, with longer retention times, diminishes the need for DACs to continuously drive the photonic devices, leading to further power savings as the processor scales (Figure 1c, inset). As the DEOAM capacitance decreases, further power savings can be realized up to the limit of the junction capacitance of the PN junction (<< 50 fF [li_25gbs_2011, li_3-d-integrated_2021]). Hence, integrating DEOAM and reducing the reliance on DACs paves the way for more efficient computing.
@@
In neuromorphic processors, analog memory is typically situated close to or directly integrated with the computing unit to minimize data movement, thereby reducing energy consumption and latency [sebastian_memory_2020, tsai_recent_2018]. Data movement in memory accesses can consume more than 10 times the energy of multiplication and addition operations [horowitz_11_2014]. Recent demonstrations of in-memory compute, such as Mythic’s Analog Matrix Processor or NeuRRAM, achieve up to tens of TOPS/W in compute efficiency by co-locating the analog memory with the compute unit [fick_mythic_2018, wan_compute--memory_2022, feldmann_parallel_2021, ambrogio_analog-ai_2023]. This highlights DEOAM’s close integration with the MRR, which acts as the compute unit in the optical domain, thereby minimizing data movement and enhancing compute efficiency.
@@ 2.1  Electro-Optic Characteristics
The neuromorphic photonic hardware with integrated DEOAM is shown in Figure 2. The MRRs are PN junction doped modulators operating in depletion mode (reverse bias) and are connected to analog memory cells consisting of 100x100 $\mu m$ metal-oxide-metal capacitors that use interdigitated fingers to increase capacitance. The MRRs have different ring perimeters starting with 144.248 $\mu m$ and 50 $\mu m$ PN junctions, increasing by 50 nm to avoid resonance collision. Key figures of merit for DEOAM are summarized in Table 1, and measurement results are shown in Figure 3. Figure 2a and b show the DEOAM circuit and its operation. Weights are loaded as voltages on the capacitive memory, affecting the MRR’s resonance. The capacitive memory holds the voltage on the MRR until the charge leaks away, returning the MRR to its resonance. Figure 2c shows the chip and measurement setup for static optical and time domain measurements. Free spectral range, inter-channel spacing, Q-factor, and extinction ratio are crucial optical characteristics that determine the maximum number of MRRs and weights in a weight bank in a specific wavelength range. Therefore, if additional weights are needed in each row of a weight bank, solutions like adding more cores, reusing hardware, or employing time multiplexing may be needed for processing neural network data. In a neuromorphic photonic weight bank, light modulation is initiated at the input data modulators. The light then travels through the weight banks before being absorbed by photodetectors (PD) and amplified by transimpedance amplifiers (TIA). The group index helps characterize the group delay, which determines the overall compute time when combined with the delay from the PD and TIA. Since the delay from the PD and TIA depends on their design and technology, we report compute time based primarily on the group delay. State-of-the-art demonstrations of PDs and TIAs yield latencies ranging in tens of ps [atef_optical_2013, babar_83-ghz_2022, awny_40_2015, kudszus_46-ghz_2003, awny_235_2016]. The wavelength and modulation efficiencies are also important for characterizing the tunability and range of weights for the neural network.
@@
Figure 2: The architecture and operation of the neuromorphic photonic prototype with integrated DEOAM. a) The analog memory circuit consists of a 1x4 MRR weight bank with each MRR connected to a capacitive analog memory. Transmission gates are used to enable/disable (EN) writing and switches are used to reset (RST) the analog memories. Laser light is injected into grating couplers through the IN port and is measured at the THRU and DROP ports. b) The operation of the DEOAM circuit is shown. The PN junction MRR is modeled as a reversed biased diode, and the transmission gate is modeled as an ideal switch. At time 1, the laser has been tuned to the MRR’s resonance, and the transmission gate is connected such that the voltage source (VIN) can drive the analog memory (VPN). At time 2, voltage is driven onto the capacitor and MRR, causing the MRR’s resonance to shift. At time 3, the transmission gate disconnects the voltage source from the capacitor, and the capacitor holds the charge on the MRR until it leaks away, returning the MRR resonance back to its original state. c) The chip is wire bonded to a custom printed circuit board with thermal control and high speed electrical lines. Details about the hardware operation and measurement setup are expanded in Methods.
@@
Key FoMs to characterize the capacitive analog memory are retention time, write time, read time, and power consumption. Since data written to DEOAM is immediately available in the optical domain, read time is effectively zero. Retention time is dependent on leakage, and leakage is dependent on optical power and photoresponse of the PN junction MRR [yu_using_2012]. Figures 3a and b reveal that the retention time of 0.8345 ms for one time constant can be extended if lower optical power is used since lower leakage is observed, but this involves a tradeoff between retention time and signal-to-noise ratio. Energy consumption, estimated to be 55.97 pJ/write, is linked to capacitor leakage and charge replenishment to maintain weight values. Figures 3c, d, and e show that the write time is about 40 to 50 ns for one time constant, enabling MHz write speeds. From Figure 3, the analog memory bit precision is about 5 bits and can be extracted from the time domain responses based on $\log_{2}(\frac{\overline{\mu}_{\mathrm{max}}-\overline{\mu}_{\mathrm{min}}}{\sigma})=N_{\mathrm{b}}$ where $\overline{\mu}_{\mathrm{max}}$ and $\overline{\mu}_{\mathrm{min}}$ are the mean values at the maximum and minimum range of the analog memory, respectively, $\sigma$ is the standard deviation, and $N_{\mathrm{b}}$ is the bit precision of the analog memory [tait_feedback_2018].
@@
Table 1: Experimental measurement results of the electro-optic analog memory circuit. Write time and retention time measured with 0.5 dBm optical bus power in the circuit. Details regarding measurement setup and optical characteristics can be found in Methods and Supplementary Information.
@@
Free Spectral Range FSRFSR - 4.67 - nm
Full Width Half Max. FWHMFWHM - 0.019 - nm
@@
Retention Time (10-90% Level) $t_{\mathrm{ret}}$ 0.3330 0.5735 0.7957 ms
Retention Time (One Time Constant) $\tau_{\mathrm{ret}}$ 0.2590 0.5527 0.8345 ms
Compute Time [^2] Estimated by the group delay through the circuit and resonance build-up in the MRRs. The summation of path delay ($\tau_{\mathrm{path}}=Ln_{\mathrm{g}}/c$ where $\tau_{\mathrm{path}}$ is the path delay, $L$ is the path length, and $c$ is the speed of light) and MRR resonance build-up time ($\tau_{\mathrm{MRR}}=FRn_{\mathrm{g}}/c$ where $R$ is the ring radius) defines compute time [geuzebroek_ring-resonator-based_2006]. $\tau_{\mathrm{read}}$ 70.87 - 79.51 ps
Write Time - Rising Edge (10-90% Level) $t_{\mathrm{rise}}$ 65.5 75.3 91.3 ns
Write Time - Falling Edge (10-90% Level) $t_{\mathrm{fall}}$ 76.2 89.4 121.2 ns
Write Time - Rising Edge (One Time Constant) $\tau_{\mathrm{rise}}$ 35.5 43.8 52.0 ns
Write Time - Falling Edge (One Time Constant) $\tau_{\mathrm{fall}}$ 51.0 63.3 71.7 ns
Bit Precision $N_{\mathrm{b}}$ 5.33 5.65 5.97 bits
Capacitance $C_{\mathrm{MEM}}$ - 14.005 - pF
@@
Figure 3: Electro-optic characteristics of the DEOAM. a) Leakage shows a dependence on optical power incident on the doped MRR. Error bands represent the standard deviation in the leakage. b) Retention time from leakage is 0.835 ms for one time constant ($\tau_{\mathrm{ret}}$) at 0.5 dBm optical bus power. The decay shows a nonlinear response from the ring’s Lorentzian response. c) With the transmission gate on, 2 V is written to analog memory using a 5 ns edge rate, and the resulting write time is 45.0 ns for one time constant ($\tau_{\mathrm{rise}}$). d) With a source voltage of 2 V on $V\textsubscript{IN}$ and toggling the transmission gate from off to on, write time is 40.0 ns ($\tau_{\mathrm{rise}}$). e) With the transmission gate on and writing 0 V to analog memory at an edge rate of 5 ns, write time is 51.0 ns for one time constant ($\tau_{\mathrm{fall}}$). Further discussions on the electro-optic results are expanded in Supplementary Information.